OpenStudy (anonymous):

simplify the complex fraction 5+1/4 / 2+2/3

OpenStudy (tkhunny):

I cannot understand what you have written. Please use parentheses, perhaps lots of them, to clarify intent. Maybe: \(\dfrac{5+\dfrac{1}{4}}{2 + \dfrac{2}{3}}\)?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

You could have written [5 + (1/4)]/[2 + (2/3)] Parentheses add clarity! Find the least common denominator if all denominators that appear.

OpenStudy (anonymous):

okay sorry for typing it so bad and i got 1

OpenStudy (tkhunny):

It's not a matter of typing badly. It's simply a matter of typing so that others can understand. Use parentheses! You got 1 for what? Did you do my little exercise? Find the least common denominator if all denominators that appear.

OpenStudy (anonymous):

im not sure

OpenStudy (tkhunny):

?? Have you ever added fractions? How would you add these? \(\dfrac{1}{4} + \dfrac{2}{3}\)

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

3/12

OpenStudy (tkhunny):

Where did you get 12? Why did you use that value? What happened that caused you to think that you might need 12?

OpenStudy (anonymous):

i dontknow what im doing

OpenStudy (tkhunny):

Go look at your original expression. Do NOT multiple numerator and denominator by different values. Find the least common denominator of all the denominators that exist. Numbers involved are 5, 1/4, 2, 2/3. What is the least common denominator amongst those four values?

OpenStudy (anonymous):

... no... just change the 2 complex fraction into improper fractions then remember, dividing by a fraction is the same as multiplying by its reciprocal

OpenStudy (anonymous):

\[\frac{ 5 +\frac{1}{4}}{2+\frac{2}{3}}\]\[5= \frac{20}{4}\] \[5+\frac{1}{4}=\frac{20}{4}+\frac{1}{4} = \frac{21}{4}\] \[2=\frac{6}{3}\] \[2+\frac{2}{3}=\frac{6}{3}+\frac{2}{3}=\frac{8}{3}\] \[\frac{ 5 \frac{1}{4}}{2\frac{2}{3}}= \frac{\frac{21}{4}}{\frac{8}{3}}= \frac{21}{4}*\frac{3}{8}= ...\] any questions?

OpenStudy (tkhunny):

\(\dfrac{5 + \dfrac{1}{4}}{2 + \dfrac{2}{3}}\cdot \dfrac{12}{12} = \dfrac{60 + 3}{24 + 16}\)

OpenStudy (tkhunny):

Whoops. Denominator is "24 + 8"